Wrong slides re-initialization
During 2 December 2011 an instrument operation anomaly was detected. The problem was due to a wrong re-initialization of the slides after the LOS calibration. Due to this anomaly, the FCE were wrongly calculated and all measurements were flagged as spikes by the L1 processor (see the plot below showing the anomalous orbits).
The anomaly started at 03:40 UTC (orbit 51024) and ended at 11:14 UTC (orbit 51027), when another re-initialization of the slides was planned before the daily gain calibration. During this period all MIPAS L1 spectra are corrupted and cannot be used for L2 retrieval.

Time and geo-location of number of detected spikes in channel A1 during 2 Dec 2011. The red crosses in the geo-located map show the anomalous measurements.